Consider the following cash flows:

Year Cash Flow
2 $ 21,400
3 39,400
5 57,400


Assume an interest rate of 8.2 percent per year.

If today is Year 0, what is the future value of the cash flows five years from now? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Future value            $

If today is Year 0, what is the future value of the cash flows ten years from now? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Future value            $
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.A=21400*(1.082)^3+39400*(1.082)^2+57400

=$130634.41(Approx).

b.A=21400*(1.082)^8+39400*(1.082)^7+57400*(1.082)^5

=$193728.66(Approx).